OpenStudy (anonymous):

Suppose you invest $2500 at an annual interest rate of 3% compounded continuously. How much will you have in the account after 7 years? Round the solution to the nearest dollar. The steps for this would be great, thanks!

OpenStudy (anonymous):

\[A=Pe ^{rt}\] P=2500 e=2.718 r=0.03 t=7 A=2500(2.718)[^{0.03(7)}\] A=2500(2.718)[^{0.21}\] A=2500(1.23367806) A=3084.20 So the answer is 3084 because you need to round to the nerest dollar